# 3D Diffuser Actor: Multi-task 3D Robot Manipulation with Iterative Error Feedback

### v2

@@ IV-C Evaluation in the real world
We collect 15 demonstrations per task where we record the keyposes, most of which naturally contain noise and multiple modes of human behavior. For example, we pick one of two ducks to put in the bowl, we insert the peg into one of two holes and we put one of three grapes in the bowl, as shown in Figure 4. 3D Diffuser Actor conditions on language descriptions and is trained to predict the next end-effector keypose. During inference, we utilize the BiRRT [37] planner provided by the MoveIt! ROS package [12] to reach the predicted poses. We evaluate 10 episodes for each task and report the success rate.
We show quantitative results in Table V and video results on our project webpage.
IV-D Run time
We measure the latency of our 3D Diffuser Actor on CALVIN in simulation, using an NVIDIA GeForce 2080 Ti graphic card. The wall-clock time of 3D Diffuser Actor is 600 ms. Notably, our model predicts end-effector keyposes sparsely, resulting in better efficiency than methods that predict actions densely at each time step. On CALVIN, on average, it takes 10 keyposes / 60 actions to complete each task using ground-truth trajectories. Our 3D Diffuser Actor can thus perform manipulation efficiently.
IV-E Limitations
@@
Our framework currently has the following limitations: 1. Our model conditions on 3D scene representations, which require camera calibration and depth information. 2. All tasks in RLBench and CALVIN are quasi-static. Extending our method to dynamic tasks and velocity control is a direct avenue of future work.
V Conclusion
We present 3D Diffuser Actor, a 3D robot manipulation policy with action diffusion. Our method sets a new state-of-the-art on RLBench and CALVIN by a large margin over existing 3D policies and 2D diffusion policies. We introduce important architectural innovations, such as 3D token representations of the robot’s pose estimate and 3D relative position denoising transformers that endow 3D Diffuser Actor with translation equivariance, and empirically verify their contribution to performance. We further test our model in the real world and show it can learn manipulation tasks from a handful of demonstrations. Our future work will attempt to train 3D Diffuser Actor policies in domain-randomized simulation environments at a large scale, to help them transfer from simulation to the real world.
VI Acknowledgements
This work is supported by Sony AI, NSF award No 1849287, DARPA Machine Common Sense, an Amazon faculty award, and an NSF CAREER award. The authors would like to thank Moritz Reuss for useful training tips on CALVIN; Zhou Xian for help with the real-robot experiments; Brian Yang for discussions, comments and efforts in the early development of this paper.
Appendix A Appendix
We explain our algorithms for keypose discovery on all benchmarks in Section A-A. We show the effect of maximum number of predicted keyposes on CALVIN in Section A-B. We detail the real-world and RLBench tasks and success conditions in Section A-C and Section A-D. We present a detailed version of our architecture in Section A-E. We list the hyper-parameters for each experiment in Section A-F. Lastly, we visualize the importance of a square cos variance scheduler for denoising the rotation estimate in Section A-G.
A-A Our heuristics for keypose discovery
For RLBench we use the heuristics from [31, 33]: a pose is a keypose if (1) the end-effector state changes (grasp or release) or (2) the velocity’s magnitude approaches zero (often at pre-grasp poses or a new phase of a task). For our real-world experiments we maintain the above heuristics and record pre-grasp poses as well as the poses at the beginning of each phase of a task, e.g., when the end-effector is right above an object of interest. We report the number of keyposes per real-world task in this Appendix (Section A-C). Lastly, for CALVIN we adapt the above heuristics to devise a more robust algorithm to discover keyposes. Specifically, we track end-effector state changes and significant changes of motion, i.e. both velocity and acceleration. For reference, we include our Python code for discovering keyposes in CALVIN here:
1"""Utility functions for computing keyposes."""
2import numpy as np
3from scipy.signal import argrelextrem
4
5def motion_changed(trajectories, buffer_size):
7
8    Args:
9        trajectories: a list of 1D array
10        buffer_size: an integer indicates the
11           minimum distance of waypoints
12
13    Returns:
14        key_poses: a list of integers indicates
15           the time steps when motion of the end
16           effector changes significantly.
17    """
18    # compute velocity
19    trajectories = np.stack(
20        [trajectories[0]] + trajectories, axis=0
21    )
22    velocities = (
23        trajectories[1:] - trajectories[:-1]
24    )
25
26    # compute acceleration
27    velocities = np.concatenate(
28        [velocities, [velocities[-1]]],
29        axis=0
30    )
31    accelerations = velocities[1:] - velocities[:-1]
32
33    # compute the magnitude of acceleration
34    A = np.linalg.norm(
35        accelerations[:, :3], axis=-1
36    )
37
38    # local maximas of acceleration indicates
39    # significant motion change of the end effector
40    local_max_A = argrelextrema(A, np.greater)[0]
41
42    # consider the top 20% of local maximas
43    K = int(A.shape[0] * 0.2)
44    topK = np.sort(A)[::-1][K]
45    large_A = A[local_max_A] >= topK
46    local_max_A = local_max_A[large_A].tolist()
47
48    # select waypoints sparsely
49    key_poses = [local_max_A.pop(0)]
50    for i in local_max_A:
51        if i - key_poses[-1] >= buffer_size:
52            key_poses.append(i)
53
54    return key_poses
55
56def gripper_state_changed(trajectories):
57    """Select keyposes where the end-effector
58    opens/closes.
59
60    Args:
61        trajectories: a list of 1D array
62
63    Returns:
64        key_poses: a list of integers indicates
65           the time steps when the end-effector
66           opens/closes.
67    """
68    trajectories = np.stack(
69        [trajectories[0]] + trajectories, axis=0
70    )
71    openess = trajectories[:, -1]
72    changed = openess[:-1] != openess[1:]
73
74    key_poses = np.where(changed)[0].tolist()
75
76    return key_poses
77
78def keypoint_discovery(trajectories, buffer_size=5):
80
81    Args:
82        trajectories: a list of 1D array
83        buffer_size: an integer indicates the
84           minimum distance of waypoints
85
86    Returns:
87        key_poses: an Integer array indicates the
88           indices of keyposes
89    """
90    motion_changed = motion_changed(
91        trajectories, buffer_size
92    )
93
94    gripper_changed = (
95        gripper_state_changed(trajectories)
96    )
97    one_frame_before_gripper_changed = [
98        i - 1 for i in gripper_changed if i > 1
99    ]
100
101    last_frame = [len(trajectories) - 1]
102
103    key_pose_inds = (
104        moition_changed +
105        gripper_changed.tolist() +
106        one_frame_before_gripper_changed.tolist() +
107        last_frame
108    )
109    key_pose_inds = np.unique(key_pose_inds)
110
111    return key_pose_inds
Figure 5: Zero-shot long-horizon performance on CALVIN with varying number of keyposes allowed at test time.
Refer to caption Refer to caption
A-B Evaluation on CALVIN with varying number of maximum keyposes allowed at test time
We show zero-shot long-horizon performance on CALVIN with varying number of keyposes allowed at test time in Figure 5. The performance saturates with more than 300 keyposes.
A-C Real-world tasks
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
  6. 6.
  7. 7.
  8. 8.
  9. 9.
  10. 10.
  11. 11.
  12. 12.
A-D RLBench tasks under PerAct’s setup
We provide an explanation of the RLBench tasks and their success conditions under the PerAct setup for self-completeness. All tasks vary the object pose, appearance and semantics, which are not described in the descriptions below. For more details, please refer to the PerAct paper [69].
@@ A-D RLBench tasks under PerAct’s setup
A-E Detailed Model Diagram

### v3

@@ 4.3 Evaluation in the real world
We validate 3D Diffuser Actor in learning manipulation tasks from real-world demonstrations across 12 tasks. We use a Franka Emika robot equipped with a Azure Kinect RGB-D sensor at a front view. Images are originally captured at $1280\times 720$ resolution and downsampled to a resolution of $256\times 256$. During inference, we utilize the BiRRT [kuffner2000rrt] planner provided by the MoveIt! ROS package [coleman2014reducing] to reach the predicted poses. We collect 15 demonstrations per task, most of which naturally contain noise and multiple modes of human behavior. For example, we pick one of two ducks to put in the bowl, we insert the peg into one of two holes and we put one of three grapes in the bowl. We evaluate 10 episodes for each task and report the success rate. We show quantitative results in Table 5 and video results on our project webpage. 3D Diffuser Actor effectively learns real-world manipulation from a handful of demonstrations.
Runtime We compare the control frequency of our model against ChainedDiffuser [xian2023chaineddiffuser] and 3D Diffusion Policy [Ze2024DP3] on CALVIN, using an NVIDIA 2080 Ti graphic card. The inference speed of 3D Diffuser Actor, ChainedDiffuser and 3D Diffusion Policy is 600ms, 1170ms (50 for keypose detection and 1120 for trajectory optimization) and 581ms for executing 6, 6 and 4 end-effector poses. The respective control frequency is 10 Hz, 5.1 Hz and 5.2 Hz.
Limitations
@@
Despite its SOTA performance with large margins over existing methods, our framework currently has the following limitations: 1. It requires camera calibration and depth information, same as all 3D policies. 2. All tasks in RLBench and CALVIN are quasi-static. Extending our method to dynamic tasks and velocity control is a direct avenue of future work. 3. It is on average slower than non-diffusion policies. This can be improved by recent techniques on reducing the inference steps of diffusion models.
Please, refer to our Appendix for more experiments and details: Section  A.1 for robustness of 3D Diffuser Actor to depth noise, Section A.2 for discussion of failure cases, Section A.3 and  A.4 for descriptions of tasks in RLBench and the real world, Section A.5 for descriptions of baselines, Section A.5.1 for implementation details of re-training 3D Diffusion Policy on CALVIN, Section A.6 for keypose discovery on CALVIN. Video results can be found in our supplementary file.
5 Conclusion
We present 3D Diffuser Actor, a manipulation policy that combines 3D scene representations and action diffusion. Our method sets a new state-of-the-art on RLBench and CALVIN by a large margin over existing 2D and 3D policies, and learns robot control in the real world from a handful of demonstrations. 3D Diffuser Actor builds upon recent progress on 3D tokenized scene representations for robotics and on generative models and shows how their combination is a powerful learning from demonstration method. Our future work includes learning 3D Diffuser Actor policies from suboptimal demonstrations and scaling up training data in simulation and in the real world.
[…]
Appendix A Additional Experimental Results and Details
A.1 Robustness to noisy depth information on RLBench
We evaluate the robustness of 3D Diffuser Actor to noisy depth information. Our results are presented in Table 6. We adopt the single-view setup on RLBench, reporting success rates on 10 tasks using the front camera view. We add Gaussian noise with a variance of $0.01$ and $0.03$ to the 3D location of each pixel. We evaluate the final checkpoints across 3 seeds with 25 episodes for each task. With mild perturbation, 3D Diffuser Actor can still predict actions precisely, achieving an average success rate of $72.4\%$. We only observed $6\%$ absolute performance drop. With strong perturbation, 3D Diffuser Actor obtains an average success rate of $50.1\%$. Even though 3D Diffuser Actor was trained with clean depth maps, it performs reasonably well with noisy depth information.
Avg. close open sweep to meat off turn slide put in drag push stack $\sigma$ Success. jar drawer dustpan grill tap block drawer stick buttons blocks 0 (clean) 78.4 82.7±1.9 89.3±7.5 94.7±1.9 88.0±5.7 80.0±8.6 92.0±0.0 77.3±3.8 98.7±1.9 69.3±5.0 12.0±3.7 0.01 72.4 80.0±3.3 81.3±1.9 68.0±3.3 85.3±5.0 82.7±7.5 92.0±0.0 65.3±3.8 93.3±3.8 70.7±1.9 5.3±1.9 0.03 50.1 41.3±6.8 25.3±10.0 45.3±3.8 74.7±3.8 89.3±3.8 89.3±3.8 18.7±1.9 48.0±3.3 68.0±5.7 1.3±1.9 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] $\sigma=0$ (clean) $\sigma=0.01$ $\sigma=0.03$
Table 6: Multi-Task performance on noisy depth information. Top: We evaluate the robustness of 3D Diffuser Actor to noisy depth information. We adopt the single-view setup in RLBench, reporting success rates on 10 tasks using the front camera view. We add Gaussian noise with a variance of $0.01$ and $0.03$ to the 3D location of each pixel. We evaluate the final checkpoints across 3 seeds with 25 episodes for each task. 3D Diffuser Actor performs reasonably well with noisy depth information. Bottom: Visualization of the 3D location of each pixel with different levels of noise.
A.2 Failure cases on RLBench
We analyze the failure modes of 3D Diffuser Actor on RLBench. We categorize the failure cases into 4 types: 1) pose precision, where predicted end-effector poses are too imprecise to satisfy the success condition, 2) instruction understanding, where the policy fails to understand the language instruction to grasp the target object, 3) long-horizon task completion, where the policy fails to complete an intermediate task of a long-horizon task, and 4) path planning, where the motion planner fails to find an optimal path to reach the target end-effector pose.
We conduct the experiment with single-view setup on RLBench, aggregating the failure cases among 10 manipulation tasks. As shown in Fig. 2, the major failure mode is imprecise prediction of end-effector poses. This failure mode often occurs in manipulation tasks that require high precision of predicted poses, such as stack blocks, open drawer and turn tap. Confusion of language instruction is another major failure mode. This mode often occurs in the scene with multiple objects, where the policy fails to grasp the target object specified in the instruction, such as stack blocks, close jar and push buttons. On RLBench, we deploy a sample-based motion planner to find the path to reach the keypose. The motion planner could sometimes fail to find the path. Lastly, 3D Diffuser Actor might fail to complete an intermediate task, such as open drawer of put item in drawer task.
Figure 2: Failure cases on RLBench on the setup of GNFactor. We categorize the failure cases into 4 types: 1) precise pose prediction, where predicted end-effector poses are too imprecise to satisfy the success condition, 2) instruction understanding, where the policy fails to understand the language instruction to grasp the target object, 3) long-horizon task completion, where the policy fails to complete an intermediate task of a long-horizon task, and 4) path planning, where the motion planner fails to find an optimal path to reach the target end-effector pose. Imprecise action prediction and confusion of language instruction are two major failure modes on RLBench.
A.3 RLBench tasks under multi-view setup
We provide an explanation of the RLBench tasks and their success conditions under the multi-view setup for self-completeness. All tasks vary the object pose, appearance and semantics, which are not described in the descriptions below. For more details, please refer to the PerAct paper [shridhar2023perceiver].
@@ A.3 RLBench tasks under multi-view setup
A.4 Real-world tasks
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
  6. 6.
  7. 7.
  8. 8.
  9. 9.
  10. 10.
  11. 11.
  12. 12.
A.5 Additional details on baselines
We provide more details of the baselines used in this paper. For RLBench, we consider the following baselines:
  1. 1.
    C2F-ARM-BC [james2022coarse], a 3D policy that iteratively voxelizes RGB-D images and predicts actions in a coarse-to-fine manner. Q-values are estimated within each voxel and the translation action is determined by the centroid of the voxel with the maximal Q-values.
  2. 2.
    PerAct [shridhar2023perceiver], a 3D policy that voxelizes the workspace and detects the next voxel action through global self-attention.
  3. 3.
    Hiveformer [guhur2023instruction], a 3D policy that enables attention between features of different history time steps.
  4. 4.
    PolarNet [Chen2023PolarNet3P], a 3D policy that computes dense point representations for the robot workspace using a PointNext backbone [Qian2022PointNeXtRP].
  5. 5.
    RVT [goyal2023rvt], a 3D policy that deploys a multi-view transformer to predict actions and fuses those across views by back-projecting to 3D.
  6. 6.
    Act3D  [gervet2023act3d], a 3D policy that featurizes the robot’s 3D workspace using coarse-to-fine sampling and featurization. We observed that Act3D does not follow the same setup as PerAct on RLBench. Specifically, Act3D uses different 1) 3D object models, 2) success conditions, 3) training/test episodes and 4) maximum numbers of keyposes during evaluation. For fair comparison, we retrain and test Act3D on the same setup.
  7. 7.
    GNFactor [ze2023gnfactor], a 3D policy that co-optimizes a neural field for reconstructing the 3D voxels of the input scene and a PerAct module for predicting actions based on voxel representations.
We consider the following baselines for CALVIN:
  1. 1.
    MCIL [lynch2020language], a multi-modal goal-conditioned 2D policy that maps three types of goals–goal images, language instructions and task labels–to a shared latent feature space, and conditions on such latent goals to predict actions.
  2. 2.
    HULC [mees2022hulc], a 2D policy that uses a variational autoencoder to sample a latent plan based on the current observation and task description, then conditions on this latent to predict actions.
  3. 3.
    RT-1 [brohan2022rt], a 2D transformer-based policy that encodes the image and language into a sequence of tokens and employs a Transformer-based architecture that contextualizes these tokens and predicts the arm movement or terminates the episode.
  4. 4.
    RoboFlamingo [li2023vision], a 2D policy that adapts existing vision-language models, which are pre-trained for solving vision and language tasks, to robot control. It uses frozen vision and language foundational models and learns a cross-attention between language and visual features, as well as a recurrent policy that predicts the low-level actions conditioned on the language latents.
  5. 5.
    SuSIE [black2023zero], a 2D policy that deploys an large-scale pre-trained image generative model [brooks2023instructpix2pix] to synthesize visual subgoals based on the current observation and language instruction. Actions are then predicted by a low-level goal-conditioned 2D diffusion policy that models inverse dynamics between the current observation and the predicted subgoal image.
  6. 6.
    GR-1 [wu2023unleashing], a 2D policy that first pre-trains an autoregressive Transformer on next frame prediction, using a large-scale video corpus without action annotations. Each video frame is encoded into an 1d vector by average-pooling its visual features. Then, the same architecture is fine-tuned in-domain to predict both actions and future observations.
  7. 7.
    ChainedDiffuser [xian2023chaineddiffuser], a combination of Act3D [gervet2023act3d] that predicts end-effector poses at key frames and a 3D trajectory predictor that generates intermediate trajectories to reach target end-effector poses.
  8. 8.
    3D Diffusion Policy [Ze2024DP3], that encodes sparsely sampled point cloud into 3D representations using a PointNeXt [Qian2022PointNeXtRP] encoder. The 3D representations are average pooled into a 1D feature vector. Then, a UNet conditions on the holistic 3D scene representations and predicts the robot end-effector poses. Notably, in its original implementation, 3D Diffusion Policy is only applied to single-task experimental setup and does not condition on language instructions. For fair comparison, we update the architecture to enable language conditioning. See Section A.5.1 for more details.
A.5.1 Re-training of 3D Diffusion Policy on CALVIN
We compare 3D Diffuser Actor against 3D Diffusion Policy on CALVIN. However, in its original implementation, the 3D Diffusion Policy does not condition on language instructions and is applied only in a single-task setup. For a fair comparison, we enabled language conditioning in the 3D Diffusion Policy by updating its architecture with additional point-cloud-to-language cross-attention layers. We used a language encoder similar to 3D Diffuser Actor to encode language instructions into latent embeddings and applied three cross-attention layers among average pooled point-cloud features and all language tokens. To ensure the maximal performance, we do not use the simplified backbone of DP3 (Simple DP3), but the original DP3 architecture.
We adopt the common setup in CALVIN, using both the front and wrist camera view. Based on the suggestion of the paper [Ze2024DP3], we crop a $160\times 160$ and $68\times 68$ bounding box from the depth map of the front and wrist camera, sampling $1024$ points within each bounding box. We train 3D Diffusion Policy with a batch size of 5400 and a total epoch of 3000. Otherwise, we use the default hyper-parameters of 3D Diffusion Policy. We set the horizon of action prediction to 4, the number of executed actions to 3, and the number observed timesteps to 2. During inference, we allow the model to predict 360 times, resulting in a maximum temporal horizon of 1080 actions.
A.6 Keypose discovery
For RLBench we use the heuristics from [james2022q, james2022coarse]: a pose is a keypose if (1) the end-effector state changes (grasp or release) or (2) the velocity’s magnitude approaches zero (often at pre-grasp poses or a new phase of a task). For our real-world experiments we maintain the above heuristics and record pre-grasp poses as well as the poses at the beginning of each phase of a task, e.g., when the end-effector is right above an object of interest. We report the number of keyposes per real-world task in this Appendix (Section A.4). Lastly, for CALVIN we adapt the above heuristics to devise a more robust algorithm to discover keyposes. Specifically, we track end-effector state changes and significant changes of motion, i.e. both velocity and acceleration. For reference, we include our Python code for discovering keyposes in CALVIN here:
1"""Utility functions for computing keyposes."""
2import numpy as np
3from scipy.signal import argrelextrem
4
5def motion_changed(trajectories, buffer_size):
7
8 Args:
9 trajectories: a list of 1D array
10 buffer_size: an integer indicates the
11 minimum distance of waypoints
12
13 Returns:
14 key_poses: a list of integers indicates
15 the time steps when motion of the end
16 effector changes significantly.
17 """
18 # compute velocity
19 trajectories = np.stack(
20 [trajectories[0]] + trajectories, axis=0
21 )
22 velocities = (
23 trajectories[1:] - trajectories[:-1]
24 )
25 # compute acceleration
26 velocities = np.concatenate(
27 [velocities, [velocities[-1]]],
28 axis=0
29 )
30 accelerations = velocities[1:] - velocities[:-1]
31 # compute the magnitude of acceleration
32 A = np.linalg.norm(
33 accelerations[:, :3], axis=-1
34 )
35 # local maximas of acceleration indicates
36 # significant motion change of the end effector
37 local_max_A = argrelextrema(A, np.greater)[0]
38
39 # consider the top 20% of local maximas
40 K = int(A.shape[0] * 0.2)
41 topK = np.sort(A)[::-1][K]
42 large_A = A[local_max_A] >= topK
43 local_max_A = local_max_A[large_A].tolist()
44
45 # select waypoints sparsely
46 key_poses = [local_max_A.pop(0)]
47 for i in local_max_A:
48 if i - key_poses[-1] >= buffer_size:
49 key_poses.append(i)
50
51 return key_poses
52
53def gripper_state_changed(trajectories):
54 """Select keyposes where the end-effector
55 opens/closes.
56
57 Args:
58 trajectories: a list of 1D array
59
60 Returns:
61 key_poses: a list of integers indicates
62 the time steps when the end-effector
63 opens/closes.
64 """
65 trajectories = np.stack(
66 [trajectories[0]] + trajectories, axis=0
67 )
68 openess = trajectories[:, -1]
69 changed = openess[:-1] != openess[1:]
70 key_poses = np.where(changed)[0].tolist()
71
72 return key_poses
73
74def keypoint_discovery(trajectories, buffer_size=5):
76
77 Args:
78 trajectories: a list of 1D array
79 buffer_size: an integer indicates the
80 minimum distance of waypoints
81
82 Returns:
83 key_poses: an Integer array indicates the
84 indices of keyposes
85 """
86 motion_changed = motion_changed(
87 trajectories, buffer_size
88 )
89
90 gripper_changed = (
91 gripper_state_changed(trajectories)
92 )
93 one_frame_before_gripper_changed = [
94 i - 1 for i in gripper_changed if i > 1
95 ]
96
97 last_frame = [len(trajectories) - 1]
98
99 key_pose_inds = (
100 moition_changed +
101 gripper_changed.tolist() +
102 one_frame_before_gripper_changed.tolist() +
103 last_frame
104 )
105 key_pose_inds = np.unique(key_pose_inds)
106
107 return key_pose_inds
Figure 3: Comparison of 3D Diffuser Actor and other methods. (a) Act3D [gervet2023act3d] encodes input images into with a 2D feature extractor, lifts 2D feature maps to 3D, samples ghost points within the scene, featurizes ghost points with relative cross attention to the 3D feature cloud, and classifies each ghost point. The position is determined by the XYZ location of the ghost point with the highest score. Act3D then crops the scene with predicted end-effector position, and iterates the same classification procedure. The rotation and openess are predicted with learnable query embeddings. (b) PerAct [shridhar2023perceiver] voxelizes input images and uses a Perceiver Transformer [jaegle2021perceiver] to predict the end-effector pose. The position is determined by the XYZ location of the voxel with the highest Q-value. The rotation and openess are predicted from the max-pooled features with MLPs. (c) 3D Diffusion Policy [Ze2024DP3] encodes 3D point cloud into 1D feature vectors, followed by a A UNet that conditions on point-cloud features and denoises end-effector poses. (d) 3D Diffuser Actor uses a similar 3D scene encoder as Act3D. A relative 3D transformer conditions on 3D feature cloud to denoise end-effector poses.
Appendix B Additional Method Details
B.1 Architectural differences between our model and baselines
We describe the architectural differences between 3D Diffuser Actor and other 3D policies in Figure 3. We compare 3D Diffuser Actor with Act3D [gervet2023act3d], PerAct [shridhar2023perceiver], and 3D Diffusion Policy [Ze2024DP3]. PerAct encodes RGB-D images into voxel representations, while 3D Diffusion Policy average pools point cloud representations into a holistic 1D feature vector. Act3D encodes input images with a 2D feature extractor and lifts 2D feature maps to 3D. Act3D determines end-effector poses by iterating a coarse-to-fine classification procedure to identify the XYZ location of the ghost point with the highest classification score. Although our model adopts a similar 3D scene encoder as Act3D, 3D Diffuser Actor uses a relative 3D transformer to denoise end-effector poses.
B.2 The formulation of relative attention
To ensure translation invariance in the denoising transformer, we use the rotary positional embeddings [su2021roformer] to encode relative positional information in attention layers. The attention between query $\mathbf{q}_{i}$, key $\mathbf{k}_{j}$ and value $\mathbf{v}_{j}$ is written as:
$\mathrm{Attention}(\mathbf{q}_{i},\mathbf{k}_{j},\mathbf{v}_{j})=\frac{\exp e_{i,j}}{\sum_{l}\exp e_{i,l}}\mathbf{v}_{j}$ (1)
where $e_{i,j}=\mathbf{q}_{i}^{T}\mathbf{M}(\mathbf{p}_{j}-\mathbf{p}_{i})\mathbf{k}_{j}$, $\mathbf{p}_{i}$ / $\mathbf{p}_{j}$ denote the positions of the query / key, and $\mathbf{M}$ is a matrix function which depends only on the relative positions of points $\mathbf{p}_{i}$ and $\mathbf{p}_{j}$.
Refer to caption Refer to caption
Figure 5: Act3D architecture in more detail. Ghost points and encoded scene feature inputs are sampled based on the current end-effector position. The 3D location of ghost points are even sampled from a fixed-radius sphere, which is centered at the current end-effector position.
B.3 Detailed Model Diagram of 3D Diffuser Actor